Determine the value of c that makes the function, f(x,y) = cxy a joint probability density function over the range 0 < x < 3 and 0 < y <x.

c = ______ Round your answer to four decimal places (e.g. 98.7654).

Determine the following:

a. P (X < 1.4, Y< 2.2) = ___ Round your answer to three decimal places (e.g. 98.765).
b. P ( 1.4 X < 2.2) = ____ Round your answer to three decimal places (e.g. 98.765).
c. P (Y > 1) =___ Round your answer to three decimal places (e.g. 98.765).
D. P (X < 2, Y < 2) = ____ Round your answer to three decimal places (e.g. 98.765).
E. E(X) = ____ Round your answer to three decimal places (e.g. 98.765).
F. E(Y) = ____ Round your answer to three decimal places (e.g. 98.765).
1
Expert's answer
2020-06-15T19:21:06-0400
"f(x, y)=cxy, 0<x<3, 0<y<x"


"\displaystyle\int_{0}^3\displaystyle\int_{0}^xcxydydx=c\displaystyle\int_{0}^3x\big[{y^2\over 2}\big]\begin{matrix}
 x \\
 0
\end{matrix}dx={c\over 2}\displaystyle\int_{0}^3x^3dx="

"={c\over 8}[x^4]\begin{matrix}
 3 \\
 0
\end{matrix}={81\over 8}c=1"

Then "c=\dfrac{8}{81}\approx0.0988."

The joint probability density function then becomes


"f(x, y)=0.0988xy, 0<x<3, 0<y<x"

Determine the following:

a.

"P(X<1.4, Y<2.2)=\displaystyle\int_{0}^{1.4}\displaystyle\int_{0}^{x}0.0988xydydx="

"=0.0494\displaystyle\int_{0}^{1.4}x[y^2]\begin{matrix}
 x\\
 0
\end{matrix}dx=0.01235[x^4]\begin{matrix}
 1.4 \\
 0
\end{matrix}\approx0.047"

b.


"P(1.4<X<2.2)=P(1.4<X<2.2, 0<Y<x)="


"=\displaystyle\int_{1.4}^{2.2}\displaystyle\int_{0}^{x}0.0988xydydx=0.0494\displaystyle\int_{1.4}^{2.2}x[y^2]\begin{matrix}
 x \\
 0
\end{matrix}dx="


"=0.0494\displaystyle\int_{1.4}^{2.2}x^3dx=0.01235[x^4]\begin{matrix}
 2.2 \\
 1.4
\end{matrix}\approx0.242"

c.


"P(Y>1)=P(1<X<3, 1<Y<x)="

"=\displaystyle\int_{1}^{3}\displaystyle\int_{1}^{x}0.0988xydydx=0.0494\displaystyle\int_{1}^{3}x[y^2]\begin{matrix}
 x \\
 1
\end{matrix}dx="

"=0.0494\displaystyle\int_{1}^{3}(x^3-x)dx=0.0494[0.25x^4-0.5x^2]\begin{matrix}
 3 \\
 1
\end{matrix}\approx"

"\approx0.790"

d.


"P(X<2, Y<2)=\displaystyle\int_{0}^{2}\displaystyle\int_{0}^{x}0.0988xydydx="

"=0.0494\displaystyle\int_{0}^{2}x[y^2]\begin{matrix}
 x\\
 0
\end{matrix}dx=0.01235[x^4]\begin{matrix}
 2\\
 0
\end{matrix}\approx0.198"

e.


"E(X)=\displaystyle\int_{0}^{3}\displaystyle\int_{0}^{x}x\cdot0.0988xydydx="

"=0.0494\displaystyle\int_{0}^{3}x^2[y^2]\begin{matrix}
 x\\
 0
\end{matrix}dx=0.00988[x^5]\begin{matrix}
 3 \\
 0
\end{matrix}\approx2.401"

f.


"E(Y)=\displaystyle\int_{0}^{3}\displaystyle\int_{0}^{x}y\cdot0.0988xydydx="

"={0.0988\over 3}\displaystyle\int_{0}^{3}x[y^3]\begin{matrix}
 x\\
 0
\end{matrix}dx={0.01976\over 3}[x^5]\begin{matrix}
 3 \\
 0
\end{matrix}\approx1.601"